Mary's age is 2/3 that of Peter's. Two years ago Mary's age was 1/2 of what Peter's age will be in 5 years' time. How old is Peter now

Answer by TimothyLamb(4379)

x = mary age now
y = peter age now
---
x = (2/3)y
x - 2 = (1/2)(y + 5)
---
x - 0.666666666y = 0
x - 2 = (1/2)y + 5/2
---
x - 0.666666666y = 0
x - (1/2)y = 5/2 + 2
---
x - 0.666666666y = 0
x - 0.5y = 4.5
---
put the system of linear equations into standard form
---
x - 0.666666666y = 0
x - 0.5y = 4.5
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = mary age now = 18
y = peter age now = 27
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---